OpenStudy (anonymous):

How would I write (8a^-9)^-2/3 in simplest form ?

OpenStudy (anonymous):

\[a^-2 = 1/a^2 \]

OpenStudy (anonymous):

Thank you, can I just ask you how you did that? I'm not quite sure how I'm supposed to arrive at this answer.

OpenStudy (anonymous):

ok this is just a law that you have to remember. Couple of laws to remember \[a^-2 = 1/a^2\] \[(a^2)^2= a^4\] \[a ^{1/3}= \sqrt[3]{a}\] \[\sqrt{a}= a ^{1/2}\]

OpenStudy (ajprincess):